A box of books is initially at rest a distance D = 0.493 m from the end of a wooden board. The coefficient of static friction between the box and the board is μs = 0.357, and the coefficient of kinetic friction is μk = 0.280. The angle of the board is increased slowly, until the box just begins to slide; then the board is held at this angle. Find the speed of the box as it reaches the end of the board.
Answer #1

Here ,

let the angle for which the sliding starts is theta

Now , as tan(theta) = us

tan(theta) = 0.357

theta = 19.6 degree

Now , for theta = 19.6 degree

acceleration of the box , a = g * sin(theta) - uk * g * cos(theta)

a = 9.8 *(sin(19.6) - 0.280 * cos(19.6))

a = 0.711 m/s^2

distance , D = 0.493 m

let the final speed as the end is v

Using third equation of motion

v^2 - u^2 = 2 * a * D

v^2 = 2 * 0.711 * 0.93

v = 0.691 m/s

the speed of the box as it reaches the end of the board is 0.691 m/s